Ratio of surface areas of two cubes is 1:4. find the ratio of their volumes.
Answer by stanbon(48510)

Ratio of surface areas of two cubes is 1:4. find the ratio of their volumes.
--------------------------------------
Surface area of a cube = 6s^2 where s is the length of a side
-----------------------
Let the smaller surface area be 6s^2, where s=1
Then the larger surface area is 4(6s^2) = 6(4s^2) = 6(2s)^2, where side = 2
-----------------------
So, smaller cube has volume s^3
and larger cube has volumne (2s)^3 = 8s^3
-----------
Ratio of their volumnes is 1:8
======================================
Cheers,
Stan H.